OpenStudy (anonymous):

200,40,8,1 3/5 in explicit form

Directrix (directrix):

200 * 1/5 = 40 40* 1/5 = 8 8* 1/5 = 8/5 The series is geometric with common ratio 1/5. The nth term for a geometric sequence is a_n = a_1 *(1/5) ^ (n-1) where a_1 = 200. a_n = 200* (1/5) ^ (n-1) which is the explicit or nth term. If you burnish that nth term with some algebra, you will see that it can also be written as a_n = 8 * 5 ^(3 - n) . @banegirl84